# Cross-Modality Embedding of Force and Language for Natural Human-Robot Communication

### v1

@@ VII Limitations and Discussion
Shared Representation Direct mappings between forces and language through classification based methods or rule-based engines fail to capture the rich force-linguistic relationships. The learned embedding space supports generalization to novel combinations through semantic patterns. For example, after training on “gently left” and “quickly up”, the framework generates appropriate descriptions like “slightly up” for a previously unseen slow upward force by leveraging the learned semantic structure.
@@
Language Structure Scope We decomposed our language structure into <direction> and <modifier> elements (e.g., “gently left”), mapping naturally to the physical properties of force profiles - <magnitude>, <direction>, and <duration> in XYZ space. While this structured decomposition enabled a force-language associations, it does represent a simplified subset of possible force descriptions. It may not capture the full richness of natural language force descriptions. This could be improved to richer linguistic structures in future work.
@@ VIII Conclusion
We presented a framework that embeds physical force profiles and verbal descriptions in a shared latent space, enabling natural bidirectional translation between how humans describe forces in language and how robots produces them. Our dual autoencoder architecture outperformed baseline approaches by 20-30% across key metrics through its unified representation of these distinct modalities. While GloVe word embeddings showed superior force reconstruction, binary embeded encodings achieved better linguistic precision, highlighting the inherent trade-offs in representing force-language relationships. The framework demonstrated robust generalization to unseen force profiles and phrases, validating its potential for real-world human-robot interaction tasks. This work provides a foundation for more intuitive physical human-robot collaboration, particularly in applications like rehabilitation therapy where coordinated force control and verbal communication are essential.

### v2

@@ VII Limitations and Discussion
Although, our framework demonstrated effective force to language mapping, few design choices and limitations must be discussed.
@@
Shared Representation Direct mappings between forces and language through classification based methods or rule-based engines fail to capture the rich force-linguistic relationships. The shared representation through a unified latent space would enable bidirectional translation where similar forces and their linguistic descriptions can be mapped close together while preserving temporal dynamics. The learned embedding space supports generalization to novel combinations through semantic patterns. For example, after training on “gently left” and “quickly up”, the framework generates appropriate descriptions like “slightly up” for a previously unseen slow upward force by leveraging the learned semantic structure.
@@
Language Structure Scope We decomposed our language structure into $<$direction$>$ and $<$modifier$>$ elements (e.g., “gently left”), mapping naturally to the physical properties of force profiles - $<$magnitude$>$, $<$direction$>$, and $<$duration$>$ in XYZ space. While this structured decomposition enabled a comprehensible force-language associations, it does represent a simplified subset of possible force descriptions. It may not capture the full richness of natural language force descriptions. This could be improved to richer linguistic structures in future work.
@@ VIII Conclusion
We presented a framework that embeds physical force profiles and verbal descriptions in a shared latent space, enabling natural bidirectional translation between how humans describe forces in language and how robots produces them. Our dual autoencoder architecture outperformed baseline approaches by 20-30% across key metrics through its unified representation of these distinct modalities. While S-BERT word embeddings showed superior force reconstruction, binary embeded encodings achieved better linguistic precision, highlighting the inherent trade-offs in representing force-language relationships. The framework demonstrated robust generalization to unseen force profiles and phrases, validating its potential for real-world human-robot interaction tasks. This work provides a foundation for more intuitive physical human-robot collaboration, particularly in applications like rehabilitation therapy where coordinated force control and verbal communication are essential.